Find the general solution of the second order constant coefficient linear ODEs7. Find the general solution of the second order constant coefficient linear ODE. (a) y +2y = 0 (b) 2y – 3y +y=0 (c) y – 2
Answer #1

a) y +2y = 0 It quxillan equation is, m²+2m = 0 => mcm+2)=0 hab m = 0, -2. As roots are real & distinct, General solution isc) The y-27-240 aurollary equation is, m²-2m- 2 = 0. m = -(-2) + √(-2) 2_401) (-2) m ayo C 2017 Cts Ctmm m = 27 J13 2 m = 2e The yllt 24 - 84 = 0 auxillary equation is, Door m²+2m - 8 = 0 m² +4 m-am - 8 = 0 mcmt4) -2 (m+4)=ons (m+4) (m-2) ao m = -g y-40+ 4y =o. The ausullay equation A, m²-4m +4=0 ² (m. 2)² = 0 => m= 2,2 As roots are real & repeated, solution is, y = (